OpenStudy (anonymous):

Please help me! If 5p = -100 + 3Q^2, which of the following could Q possibly be?

OpenStudy (anonymous):

Q = square root of 5 times p minus 100 over 3 Q = square root of 5 times p plus 100 over 3 Q = square root of negative 5 times p plus 100 over 3 Q = square root of negative 5 times p minus 100 over 3

OpenStudy (anonymous):

Is it 3Q^2??

OpenStudy (anonymous):

Yes, my bad

OpenStudy (anonymous):

Could you use the equation button to rewrite the whole thing? Im having trouble reading it.

OpenStudy (anonymous):

\[5p = -100 + 3Q^2\]

OpenStudy (anonymous):

Give me a sec. :)

OpenStudy (anonymous):

Those are you answer choices?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@DW101796 ?

OpenStudy (anonymous):

Sorry, was helping someone else out. Im getting something completely different. One sec.

OpenStudy (anonymous):

Heres what I'm getting. \[Q = \sqrt(5/3) \sqrt(p+20)\]

OpenStudy (anonymous):

Are you sure you typed the problem correct?

OpenStudy (anonymous):

Yes. I typed it exactly as it was.

OpenStudy (precal):

its the second one

OpenStudy (precal):

\[5p=-100+3Q^2\] \[5p+100=3Q^2\] \[\frac{ 5p+100 }{ 3 }=Q^2\] \[\sqrt{\frac{ 5p+100 }{ 3}}=Q\]